Consider the shuttle carrying the astronauts is dropped from 100ft. The position 𝑠 of the shuttle is given by the following function as a function of time: 𝑠(𝑑) = βˆ’16𝑑 2 + 100.s. What is the domain of this function?


1
Expert's answer
2021-09-27T02:31:45-0400

"0\le s(t)\le 100"


For "s(t)=0" :

"-16t^2+100=0"

"t=10/4=2.5\ s"


For "s(t)=100" :

"-16t^2+100=100"

"t=0"


Domain:

"0\le t \le 2.5"